OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and that passes through the given point. y = 2x + 7; (3, 11)

jimthompson5910 (jim_thompson5910):

same idea as last time y = 2x+7 has a slope of 2 So any parallel line to this one also has a slope of 2, therefore m = 2 The point is (3,11) so (x,y) = (3,11) ---> x = 3, y = 11

jimthompson5910 (jim_thompson5910):

y = mx+b y = 2x+b 11 = 2*3 + b b = ???

OpenStudy (anonymous):

i multiplied 2*3 and got 6 of course now this is how it looks to me

jimthompson5910 (jim_thompson5910):

11 = 2*3 + b 11 = 6 + b b = ???

OpenStudy (anonymous):

\[11=6\]

OpenStudy (anonymous):

do i have to divide something on both sides ??

jimthompson5910 (jim_thompson5910):

6 plus NUMBER = 11 what must replace NUMBER ?

OpenStudy (anonymous):

4 >.>

jimthompson5910 (jim_thompson5910):

6 + 4 = 11 ?

OpenStudy (anonymous):

oops 5

jimthompson5910 (jim_thompson5910):

6 + 5 = 11, yep

jimthompson5910 (jim_thompson5910):

so b = 5

jimthompson5910 (jim_thompson5910):